### v1

# Rethinking Discrete Speech Representation Tokens for Accent Generation

Abstract
Discrete Speech Representation Tokens (DSRTs) have become a foundational component in speech generation. While prior work has extensively studied phonetic and speaker information in DSRTs, how accent information is encoded in DSRTs remains largely unexplored. In this paper, we present the first systematic investigation of accent information in DSRTs. We propose a unified evaluation framework that measures both accessibility of accent information via a novel Accent ABX task and recoverability via cross-accent Voice Conversion (VC) resynthesis. Using this framework, we analyse DSRTs derived from a variety of speech encoders. Our results reveal that accent information is substantially reduced when ASR supervision is used to fine-tune the encoder, but cannot be effectively disentangled from phonetic and speaker information through naive codebook size reduction. Based on these findings, we propose new content-only and content-accent DSRTs that significantly outperform existing designs in controllable accent generation. Our work highlights the importance of accent-aware evaluation and provides practical guidance for designing DSRTs for accent-controlled speech generation.
Rethinking Discrete Speech Representation Tokens for Accent Generation
Jinzuomu Zhong1thanks: Equal contribution., Yi Wang [^1] footnotemark: 1, Korin Richmond1, Peter Bell1, 1Centre for Speech Technology Research, University of Edinburgh Correspondence: {jinzuomu.zhong, korin.richmond}@ed.ac.uk
1 Introduction
Inspired by the tokenisation schemes employed by text-based Large Language Models (LLMs), discrete speech tokenisation has emerged as an active research area for bridging speech and LLMs. By compressing speech signals or learned continuous representations into a finite set of discrete units, discrete speech tokens enable efficient storage of complex acoustic information while remaining inherently compatible with LLM architectures. As a result, they are rapidly becoming established as a foundational representation across a wide range of speech tasks (Guo et al., 2025).
Discrete speech representation tokens (DSRTs), quantised from speech representations, have consequently seen huge success in various speech generation tasks, including Speech Language Models (SpeechLMs) (Cui et al., 2025), Zero-Shot Text-to-Speech (ZS-TTS) (Kharitonov et al., 2023), Speech-to-Speech Translation (S2ST) (Lee et al., 2022) and full-duplex dialogue systems (Nguyen et al., 2023).
Despite the significant progress brought by DSRTs, the representation of a speaker’s accent is largely overlooked in the design, evaluation, and application of the tokens. Prior user perception studies have demonstrated a similarity–attraction effect, whereby listeners prefer accents similar to their own in spoken interaction (Dahlbäck et al., 2007). Unfortunately, ZS-TTS systems are shown to hallucinate accents that differ from those of the reference speaker (Zhong et al., 2025b), while spoken dialogue systems still lack broad adoption of benchmarks or evaluation settings that consider appropriate, accent-controlled speech generation (Cheng et al., 2025).
Although some recent zero-shot TTS (ZS-TTS) systems have demonstrated some ability to mimic or control the accent of generated speech using discrete speech representation tokens (DSRTs), how much accent information is encoded in these tokens remains unexplored (Du et al., 2024; Zhang et al., 2025; Wang et al., 2025). Existing claims – such as that naive codebook size adjustment (Zhang et al., 2025) or ASR supervision (Du et al., 2024) can facilitate controlled accent generation – lack systematic investigation. To this end, we ask the following research questions: (1) How do different design choices in DSRTs influence the amount of accent information they encode? (2) How can these insights be leveraged to enable more controllable accent generation, e.g. in voice conversion (VC)?
Existing DSRT evaluation frameworks focus on phonetic and speaker information and rarely consider accent (Polyak et al., 2021). We extend the investigation of accessibility to accent information, with an accent ABX method we proposed that evaluates the discriminability of representations for words uttered in different accents. Additionally, we propose to evaluate the recoverability of accent, speaker, and phonetic information by cross-accent VC, resynthesising from DSRTs from source speakers and speaker IDs from target speakers of different accents. Using the proposed framework, we observe the following key findings: (1) Accent information is most prominent in mid-early layers of HuBERT, different from speaker or phonetic information distribution. (2) Naive adjustment of codebook sizes provides only very limited disentanglement of accent, speaker and content information. (3) Predominant design of DSRTs for speech generation (quantising a later layer in a speech representation model, or using ASR supervison) discards most accent information.
Our main contributions are:
  • To the best of our knowledge, this is the first work to incorporate accent in evaluating DSRTs, and to apply ABX to a study of accent.
  • Based on our findings, we propose quantisation schemes that are appropriate for accent-preserving VC – preserving source speaker accent – and accent-adaptive VC – adapting to target speaker accent – which achieve superior performance to existing approaches.
  • Evaluation pipeline and proposed DSRTs will be open-sourced in the future.
Figure 1: Proposed pipeline for evaluating the recoverability and accessibility of accent, speaker, and phonetic information in various Discrete Speech Representation Tokens (DSRTs).
2 Related Work
2.1 Discrete Speech Tokens
Discrete speech tokens can be largely categorised into two groups: semantic tokens [^1] A severe misnomer as these tokens contain primarily phonetic information and little semantic information (Wells et al., 2022; Choi et al., 2024). and acoustic tokens. We refer to semantic tokens as Discrete Speech Representation Tokens (DSRTs) here, naming them strictly by how these tokens are designed and trained, rather than what information some researchers assume them to possess. DSRTs are tokens obtained by quantising learned speech representations, such as HuBERT (Hsu et al., 2021) and w2v-BERT (Chung et al., 2021); while acoustic tokens are obtained by quantising waveforms. Common quantisation techniques include k-means, Vector Quantisation (VQ) (van den Oord et al., 2017), Residual Vector Quantisation (RVQ) (Zeghidour et al., 2021), and Finite-State Quantisation (FSQ) (Mentzer et al., 2024), with k-means, VQ, and FSQ primarily used for DSRTs, VQ and RVQ primarily used for acoustic tokens.
2.2 Investigation of Speech Representations
In terms of of continuous speech representations, most previous work has focused on accessibility, finding that phonetic information peaks in intermediate layers, while acoustic information dominates earlier layers (Choi et al., 2024; Pasad et al., 2021). After finetuning on the ASR task, it has been observed deeper layers become dominated by task-specific information, exhibiting higher phonetic discrimination.
In the investigation of DSRTs, Wells et al. (2022) find that DSRTs obtained by k-means on hubert-base-ls960 correspond to sub-phonetic events. Yeh and Tang (2024), using their information completeness and accessibility framework, found that DSRTs obtained from hubert-base-ls960 layer 4 contain more accessible pitch and speaker information, but less accessible phone information, than layer 9. Phonetic and speaker information in both layers consistently decrease with lower bitrates of RVQ codes, refuting claims that content and speaker disentanglement can be achieved by VQ.
2.3 DSRTs in ZS-TTS and SpeechLMs
In ZS-TTS, researchers found that DSRTs provide more accessible phonetic information than acoustic tokens do, while suffering information loss such as timbre and acoustic environment (Wang et al., 2025). Motivated by such findings, many state-of-the-art (SOTA) models utilise a hierarchical approach, first predicting DSRTs from text or source speech and then predicting acoustic tokens from DSRTs (Du et al., 2024; Lee et al., 2025; Wang et al., 2025; Zhang et al., 2025). In SpeechLMs, researchers seek to combine DSRTs with acoustic tokens for both accessible and complete/recoverable information, such as SpeechTokenizer (Zhang et al., 2024) and Mimi tokeniser (Défossez et al., 2024).
When designing such DSRTs, ZS-TTS researchers have claimed that are unverified or even contradicted by speech representation investigations. Zhang et al. (2025) choose layer 18 of hubert-large-ll60k and claim that reducing the codebook size in VQ leads to natural disentanglement of speaker, style (accent and emotion by their definition) and content, with a codebook size of 32 containing only content, and a codebook size of 8,192 containing only content and style. This claim not only contradicts the findings of Yeh and Tang (2024) that VQ or k-means with 1024 codebook size still retain significant speaker information, but also lacks verification in terms of accent and emotion. Du et al. (2024) propose to use “supervised semantic token”, obtained by injecting FSQ in an internal ASR model encoder; how ASR pretraining affects the information in discrete tokens remains unclear.
2.4 Accent Control and Generation
Unfortunately, ZS-TTS speech generation systems suffer accent hallucination, whereby they generate a hallucinated synthetic accent that deviates from the reference or prompt speech (Zhong et al., 2025b). Recent systems that have been specially built for accent generation have not yet utilised DSRTs along with powerful LLM architectures (Zhong et al., 2025b; Xinyuan et al., 2025). General ZS-TTS systems are reported to have certain accent control capabilities, but there has been little or no investigation about accent information in DSRTs. Understanding how accent information is encoded could help build more inclusive ZS-TTS systems for diverse accents.
2.5 ABX Evaluation
[…]
Previous work introduced Minimal Pair ABX (MP-ABX) (Schatz et al., 2013), where instances differ only in a central phonetic unit while sharing the same phonetic context. Prior studies (Borsos et al., 2023; Wang et al., 2025; Zhang et al., 2025) have used this metric to assess the phonetic discriminability of DSRT tokens. In addition, ABX has been extended to probe speaker-related information by defining triplets that contrast speaker identity under controlled linguistic content (de Seyssel et al., 2022). Nevertheless, to the best of our knowledge, no prior work has systematically employed machine ABX to reveal accent discriminability in speech representations or DSRTs.
3 Method
Inspired by the perspective of separately measuring information completeness and accessibility Yeh and Tang (2024), we focus on evaluating DSRTs from both aspects, but in the context of accent generation. Since completeness cannot be measured by resynthesis, we introduce recoverability as a task-grounded measure of how much accent, speaker, and phonetic information encoded in DSRTs can be recovered in resynthesised speech. Together with accessibility, we propose a pipeline that evaluates DSRTs from both synthesis-facing and representation-facing perspectives.
Figure 1 provides an overview of the proposed pipeline. First, we obtain DSRTs from several commonly used speech representations using RepCodec Huang et al. (2024) with Vector Quantisation (VQ) van den Oord et al. (2017) as the quantiser (see Section 3.1). For each DSRT configuration, we then train a unit-to-speech resynthesis model using HiFiGAN Polyak et al. (2021). To assess information recoverability, we infer unit-to-speech models with DSRTs from a source speaker and speaker ID from a target speaker with a different accent, thereby conducting cross-accent VC. The generated speech is evaluated using a combination of objective metrics and subjective listening tests (see Section 3.2). Finally, to assess information accessibility, we directly probe the DSRTs using a range of ABX setups (see Section 3.3).
3.1 Discrete Speech Representation Tokens
We choose three speech representation models to obtain Discrete Speech Representation Tokens (DSRTs): HuBERT; HuBERT finetuned for ASR (HuBERT-ft); and Whisper. We include HuBERT Hsu et al. (2021) as it provides the most widely used speech representations in various speech tasks such as SpeechLM and TTS Guo et al. (2025); Cui et al. (2025). Recently, there has been a growing trend toward using ASR-based speech representations for obtaining DSRTs. Therefore, we include HuBERT-ft and Whisper as two representative ASR-based representations, with Encoder-only and Encoder-Decoder architectures, respectively.
Following RepCodec Huang et al. (2024), we use VQ-VAE van den Oord et al. (2017) to discretise the speech representations. The model consists of three modules: Encoder and Decoder, which are both convolution layers with residual paths, and a Vector Quantisation module that quantises latent representations from Encoder output into a series of discrete tokens. The model is trained to reconstruct the speech representation, with additional Exponential Moving Average (EMA) optimisation to gradually update the codebook. For model details, we refer readers to the original paper Huang et al. (2024).
3.2 Cross-Accent Voice Conversion
After extracting DSRTs, we train a unit-to-speech HiFiGAN model for each DSRT configuration, following Polyak et al. (2021). In prior work, information recoverability is typically evaluated using resynthesis or Voice Conversion (VC) on General American English datasets, with evaluation focusing primarily on intelligibility and speaker similarity, while accent is largely ignored.
We train unit-to-speech models on data covering multiple accents to assess the generalisability of DSRTs across accents. We then perform cross-accent VC in inference, by conditioning the model on DSRTs from a source speaker and a target speaker ID with a different accent. Afterwards, during evaluation, following Zhong et al. (2025a), we explicitly assess accent, speaker, and phonetic similarity in the converted speech. Utterance-level embeddings from Accent Identification (AID) and Speaker Verification (SV) models are used to calculate cosine similarity, which serves as a proxy for perceived accent and speaker similarity. Phonetic Posteriorgrams (PPGs) are extracted, aligned, and then used to calculate phonetic distances as a proxy for perceived phonetic similarity, and to some degree, intelligibility.
By analysing the similarity of the generated speech to the source and target speech along these dimensions, we can determine whether accent information is primarily derived from the source DSRTs or overridden by the target speaker identity, thereby quantifying the amount of accent information preserved in the DSRTs.
3.3 Accent, Speaker, and Phonetic ABX
We choose ABX as a model-free method to estimate the accessibility of accent information, as well as phone and speaker information in DSRTs.
In accent ABX, instances $a$ and $x$ share the same accent, while instance $b$ differs in accent. Triplets $(a,b,x)$ are constructed to share identical lexical content rather than phonetic context, as accent distinctions arise from accent-dependent realizations of the same words or lexical units, which may involve different phone sequences. We further require that $a$, $b$, and $x$ come from different speakers to avoid trivial similarity between $a$ and $x$ due to shared speaker identity, e.g.: a triplet $(a,b,x)$ consists of: $x=$ “water” spoken by a Scottish speaker1, $a=$ “water” by a Scottish speaker2, and $b=$ “water” by a Southern English speaker3.
We also adopt speaker ABX and phone ABX as contrastive evaluations. For detailed triplet selection criteria, see Appendix A.
In practice, words vary in both accent-induced pronunciation variation and speaker-specific pronunciation patterns. While standard ABX aggregates scores over all valid triplets, we adopt a word selection scheme to improve the sensitivity of accent ABX. We first select the 100 most frequent words in the training data, compute ABX scores for each (accent $A$ and $B$, word) combination, and retain the 10% with the lowest scores as the most accent-discriminative. The final accent ABX score is obtained by averaging their ABX scores on the test set. The combination selection is performed using continuous features from GenAID Zhong et al. (2025b), a supervised accent identification model fine-tuned from XLSR Babu et al. (2021), to ensure effective word selection and a fair comparison among DSRTs. See Appendix B for details.
(a) Cross-accent VC evaluation results for information recoverability.
(b) ABX evaluation results for information accessibility.
Figure 2: Accent, speaker, and phonetic information in different DSRTs across layers and representations.
All codebook sizes and code dimensions are set to 1024 for fair comparison.
(See Figure 4 in Appendix D for additional results from cross-accent VC.)
4 Experiments
4.1 Obtaining DSRTs
We extract continuous speech representations from various layers of HuBERT222hubert-large-ll60k, available here, HuBERT-ft333hubert-large-ls960-ft, available here and Whisper444whisper-meidum.en, available here. All three speech representation models share the same Encoder architecture (24 Transformer layers) and are trained on English-only data. Here, we focus on investigating how layer choice and ASR pretraining affect the information encoded in speech representations, and leave the investigation of how multilingual pretraining or larger-scale pretraining data/model to future work.
The extracted representations are then discretised using RepCodec [^5] https://github.com/mct10/RepCodec, which we train on the train-clean-100 subset of LibriSpeech [^6] https://www.openslr.org/12 following similar practice in Huang et al. (2024) and Zhang et al. (2025), with various codebook sizes. This subset contains 100 hours of speech with "accents closer to US English" Panayotov et al. (2015). RepCodec is trained for 200,000 steps with the same hyperparameters in Huang et al. (2024).
4.2 Evaluation Dataset
To evaluate information recoverability and accessibility with respect to accent, we use the VCTK corpus [^7] https://datashare.ed.ac.uk/handle/10283/3443, which provides relatively broad coverage of native English accents. Based on the provided country labels and detailed region descriptions, we manually group speakers into 13 accent regions, as some coarse country-level labels (e.g. “US”, “England”) conflate multiple distinct accent varieties.
We withhold 3–4 speakers from each accent region (50 speakers in total, with balanced gender and accent distribution) for testing. For the remaining speakers, we use 4 accent regions that have sufficient speakers (45 speakers in total) for training unit-to-speech HiFiGAN in Section 3.2 and for selecting accent-discriminative words in Section 3.3. Out of the 13 accent regions, 4 are seen in HiFiGAN training and accent ABX word selection, 5 are likely seen by RepCodec in LibriSpeech, and the remaining 4 are unseen at any stage, allowing for analysis on the generalisability of DSRTs across accents. See Appendix C for details of accent grouping and speaker splits.
Figure 3: Accent, speaker, and phonetic information recoverability in different DSRTs across codebook sizes.
(See Figure 5 in Appendix D for additional results from cross-accent VC and ABX.)
4.3 Evaluating Information Recoverability
To evaluate information recoverability, we train unit-to-speech HiFiGAN models [^8] https://github.com/facebookresearch/speech-resynthesis for 100,000 steps with the same hyperparameters as Polyak et al. (2021), using 4 accent regions (45 speakers), as described in Section 4.2. During inference, we perform cross-accent VC using source (src) DSRTs from 12 accent regions (46 speakers, excluding Southern English), together with speaker IDs from 4 Southern English speakers as target (tgt) voices. We fix the target voices to one accent region to control for target-side accent variability, enabling a clearer analysis of how accent information is preserved in the src DSRTs.
For objective evaluation of the converted speech, we adopt metrics recommended by Zhong et al. (2025a). (1) Accent similarity: We calculate cosine similarity of accent embeddings (Accent COS-SIM) extracted from GenAID [^9] https://github.com/jzmzhong/GenAID Zhong et al. (2025b). GenAID is trained with explicit speaker-accent disentanglement, preserving little speaker information in the extracted accent embeddings. (2) Speaker similarity: We calculate cosine similarity of speaker embeddings (Speaker COS-SIM) extracted from WavLM101010wavlm-base-plus-sv, available here Chen et al. (2022). These speaker embeddings encode some accent information, evidenced by improved performance in AID when transfer learned from SV models Zuluaga-Gomez et al. (2023). (3) Phonetic similarity: We calculate Jensen-Shannon distance between aligned PPGs [^11] https://github.com/interactiveaudiolab/ppgs (PPG Distance), extracted from a phone recognition model Churchwell et al. (2024). It is worth noting that PPG distance, primarily designed for pronunciation distance, is sensitive to accent information, as accent can be characterised partially as different realisations of the same phoneme. (4) Intelligibility: We calculate Word Error Rate (WER) using whisper-medium.en transcriptions, compared against ground-truth transcriptions. Despite its evident accent bias Jahan et al. (2025), WER is included here due to its broad adoption in previous DSRT evaluation frameworks to reflect how DSRTs are commonly selected.
For Accent COS-SIM, Speaker COS-SIM, and PPG distance, we use the first 24 utterances per speaker, which is an elicitation paragraph shared across speakers. In this way, we allow for similarity/distance calculation to either source speech that provides DSRTs or target speech that provides the speaker ID of the same content, removing the influence of content on these metrics. For WER, we randomly choose 24 utterances per speaker to cover diverse contents.
For visualisation, we annotate each metric with a practical lower bound (chance-level performance) and a practical upper bound (best possible performance), which anchor the visual scale and facilitate interpretation of values. For Accent COS-SIM, PPG, and WER, the lower bound measures the source speaker ground truth (GT) with respect to target speaker GT, and the upper bound measures the source speaker copy-synthesis with respect to source speaker GT. For Speaker COS-SIM, the lower bound measures similarity between source speaker GT and target speaker GT, while the upper bound measures similarity between target speaker copy-synthesis and target speaker GT.
4.4 Evaluating Information Accessibility
Information accessibility is evaluated only on the accents seen during HiFiGAN training, ensuring sufficient numbers of speakers for each accent. As described in Section 3.3, a training set is required to select high-frequency words and accent-discriminative combinations, while a separate test set is used for ABX evaluation to prevent speaker leakage. Accents outside the four seen accents contain no more than four speakers each, severely limiting the number of valid ABX triplets.
We follow the same training–test partition as in HiFiGAN experiments for simplicity. Accent ABX is conducted using this partition, whereas phone and speaker ABX are computed only on the test split, as they do not require a training phase. Test experiments used all utterances in the split, except for the first 24 utterances to avoid identical sentences across all train and test speakers, which could bias the ABX scores toward lexical-context matching instead of reflecting true accent or speaker discrimination. Phone-level alignment (phone boundaries and labels) required by phone ABX, were obtained through the Montreal Forced Aligner with the English Arpabet dictionary McAuliffe and Sonderegger (2024); Gorman et al. (2011).
We treat the accent ABX score computed on HuBERT layer 6 continuous features as an upper bound, as this layer exhibits the strongest accent-related information. Similarly, we use the phonetic and speaker ABX scores computed on HuBERT-ft layers 3 and 9 continuous features as upper bounds, respectively, since these layers demonstrate the highest phonetic and speaker recoverability in Section 5. In addition, we evaluate the accent ABX error rate of the accent-discriminative GenAID features, which achieves a low error rate of 9.87%, validating the effectiveness of the proposed metric.
Token type
Representation
& Layer
Codebook
size
A-SIM
A-SIM
S-SIM
[…]
S-SIM
[…]
Vevo Content HuBERT L18 32 0.8143 0.9369 0.6817 0.9651 0.1718 0.0824
Content-style HuBERT L18 8192 0.8923 0.8888 0.6989 0.9499 0.1427 0.0425
Proposed Content HuBERT-ft L18 256 0.8081 0.9509 0.6782 0.9667 0.1488 0.0523
Content-accent HuBERT L9 8192 0.9541 0.8227 0.7529 0.9104 0.1265 0.0344
Table 1: Results of proposed content and content-accent tokens, compared with content and content-style tokens from Vevo (Zhang et al., 2025). Differences in DSRT design choices are marked in red, with best performances marked in bold. *: For content tokens, we want little accent information recovered from DSRTs, i.e. low A-SIM (src) and high A-SIM (tar). For content-accent or content-style tokens, we want the opposite.
5 Results
5.1 Layer Choice Matters for Accent
Layer choice has a prominent impact on the recoverability and accessibility of information in HuBERT DSRTs, shown in Figure 2 (dark green curve). Notably, accent, speaker and phonetic information are distributed differently across layers. (1) Accent information recoverability, reflected by Accent COS-SIM w.r.t. source speaker, is most prominent in mid–early HuBERT layers (L6 & L9), and decreases in earlier or latter layers, indicating that accent cues emerge after low-level acoustic processing but are progressively abstracted away in higher layers. (2) Speaker information is most recoverable in early HuBERT layers (L3), indicated by Speaker COS-SIM to the source speaker, and decreases monotonically with depth. Overall, speaker information is limited, as cross-accent VC aims to generate voice similar to that of the target speaker with speaker ID conditioning. (3) Phonetic information, measured by PPG Distance to the source speech, is most complete in middle layers (L9 & L12). Both earlier and later layers show reduced phonetic completeness, reflecting a trade-off between raw acoustic detail and higher-level abstraction. (4) Information accessibility, measured by ABX error rates (see Figure 2(b)), reveals weak accent accessibility for HuBERT across layers, despite clear inter-layer variations. Accent accessibility peaks at L12, which differs from the recoverability peak at L6 or L9. This discrepancy highlights the necessity of resynthesis-based recoverability evaluation when assessing DSRTs for accent generation: while ABX accessibility captures relative trends, it cannot reliably identify optimal layers for synthesis.
5.2 ASR Supervision Removes Accent Information
ASR supervision reduces the recoverability and accessibility of accent information in DSRTs across layers, shown in Figure 2 (blue curve for HuBERT-ft, orange curve for Whisper). (1) HuBERT consistently shows higher accent recoverability and accessibility than HuBERT-ft or Whisper, particularly in latter layers, which are more affected by ASR supervision. (2) Whisper exhibits a peak in accent and phonetic information at intermediate layers (L12 & L15), different from the distribution of information in HuBERT or HuBERT-ft.
5.3 Reducing Codebook Size Provides Limited Disentanglement
We then vary codebook sizes for three selected layers: HuBERT L18 (a common choice, as in Vevo Zhang et al. (2025)), HuBERT-ft L18 (a layer with little accent information while retaining high phonetic information), and HuBERT L9 (a layer with most accent information), shown in Figure 3. (1) Focusing on accent recoverability, varying codebook sizes has a markedly smaller effect than changing representation models or layers. For each layer, increasing codebook sizes from 32 to 2048 consistently improves accent and phonetic recoverability across all inspected DSRTs. However, the recoverability gap induced by codebook sizes variation within a layer is substantially smaller than that induced by changing layers at a fixed codebook sizes. (2) Speaker and phonetic recoverability also increase with larger codebook sizes. Taking HuBERT DSRTs as an example, Accent COS-SIM drops sharply as codebook sizes decreases from 1024 to 32, while PPG distance also rises 20.4%. Despite speaker identity being partially suppressed by the injected speaker ID during resynthesis, speaker recoverability also declines. The parallel degradation across accent, speaker, and phonetic metrics suggests minimal disentanglement when reducing codebook size.
5.4 Limitations of Commonly Used DSRTs in Speech Generation
Based on the above findings, we highlight the limitations of commonly used DSRTs here. (1) Vevo’s choice of using DSTRs based on HuBERT L18 with codebook size 32 as content tokens and codebook size 8,192 as content-style tokens Zhang et al. (2025) is suboptimal. Significant accent information is likely already lost and unrecoverable in HuBERT L18 - leading to insufficient accent information encoded in content-style tokens. Since naive reduction of codebook size cannot remove accent without harming phonetic information, content tokens achieve little accent information at the cost of intelligibility in resynthesised speech. (2) ASR-supervised DSRTs, adopted by Du et al. (2024), exhibit lower peak accent and phonetic recoverability than HuBERT DSRTs. Similar to content-style tokens, ASR-supervised DSRTs cannot encode accent sufficiently.
5.5 Proposed Content and Content-Accent Tokens
For better accent generation and control, we propose content-accent and content tokens for accent-preserving and accent-adaptive VC, respectively. A comparison of proposed tokens and Vevo tokens is presented in Table 1. Our proposed content/content-accent tokens achieve better performance than Vevo’s content/content-style tokens across all metrics, enabling better control of accent in speech generation. Demos of the converted speech will be available at [Link redacted].
6 Discussion
We observed a systematic delay of the optimal accessible layer relative to the recoverability peak. As discussed in (Yeh and Tang, 2024), earlier layers retain more complete acoustic information that can be effectively extracted by a strong HiFiGAN vocoder, whereas later layers gain phonetic discriminability from ASR-related objectives,
Across both metrics, accent information follows a distinct yet intermediate trend between phonetic and speaker information. This aligns with linguistic intuition: accent is partially realised through systematic phonetic variation and partially through speaker-specific pronunciation traits.
Layer selection enables accent control precisely because accent, speaker, and phonetic information are distributed differently across layers. Consequently, a single "style layer", as suggested by the content–style token design in Vevo, is insufficient to jointly capture multiple speech attributes such as content, accent, and emotion. A more thorough investigation of more speech attributes, and how their information are represented across multiple layers and representations, is therefore necessary to support better controllable speech generation.
Finally, our findings provide potential explanations for the hallucinated accents by ZS-TTS systems Zhong et al. (2025b). Many ZS-TTS systems rely on representations (often deep or supervised layers) where accent recoverability is already attenuated, encouraging the model to guess accent patterns in inference or default to mainstream accents in training data.
7 Conclusion
In this work, we propose a framework for evaluating discrete speech representation tokens (DSRTs) from both recoverability and accessibility perspectives, in terms of accent, speaker, and phonetic information encoded, with a focus on providing guidance for accent generation and control. Based on the results using this framework on HuBERT- and Whipser-based DSRTs, we propose content and content-accent tokens, achieving more controllable accent in voice conversion (VC) than existing works.
Limitations
Although the framework we proposed guided us to develop a pair of DSRTs achieving good control over accent in VC, we acknowledge several limitations. First, we did not evaluate the proposed DSRTs in ZS-TTS settings, as training ZS-TTS systems requires substantial computational resources. Future work will extend our evaluation to ZS-TTS to assess the generalisability of our framework beyond voice conversion. Second, the current Accent ABX test shows low accent accessibility across all evaluated DSRTs. A supervised accent representation (GenAID Zhong et al. (2025b)) achieves much lower ABX error rate (9.87%), indicating the validity of the metric. More powerful probing models are needed to better reveal subtle accent information in SSL representations and DSRTs without explicit accent-discriminative training, so as to provide more precise guidance for accent applications. Better design of DSRTs is also needed to make accent information more accessible, and thus easier to learn by downstream speech generation models. Third, our experiments are limited to English SSL models and the VCTK dataset, which contains only L1 English speaker accents. We plan to extend the framework to multilingual representations and datasets involving L2 English accents and non-English languages to improve regional and language coverage. Finally, we focus on objective metrics to evaluate DSRTs and their effectiveness in VC. We haven’t included subjective evaluation and listening test, as Zhong et al. (2025a) have shown the effectiveness of the adopted objective metrics on evaluating accent generation, correlating with subjective listening tests results.
Broader Impacts and Ethical Statements
This work improves the transparency and controllability of accent information in discrete speech representation tokens (DSRTs). The proposed framework may help mitigate accent hallucination and unintended accent drift, benefiting speakers with underrepresented accents and potentially supporting applications such as speech therapy through accent- and voice-preserving speech generation. Our results further suggest that accent control can be partially achieved via informed representation and layer selection, potentially reducing reliance on additional large-scale model training.
Although this work is intended for research use, enhanced accent controllability in speech generation could be misused for voice impersonation if deployed without safeguards. If extended to real-world or unseen speakers, practical systems should incorporate mechanisms to ensure speaker approval and to detect synthesized speech. We emphasise that our framework is designed to analyse and expose limitations of DSRTs rather than to enable unrestricted accent manipulation.
This study exclusively uses existing publicly available datasets and pretrained models for research purposes, without collecting new data or releasing derived datasets. All data and pre-trained models are used under the following licenses: Creative Commons BY 4.0 License, Creative Commons BY-NC-ND 4.0 License, Apache License 2.0, and MIT License.
Acknowledgments
This work was supported in part by the UKRI AI Centre for Doctoral Training in Responsible and Trustworthy in-the-world Natural Language Processing (Grant EP/Y030656/1), UKRI Centre for Doctoral Training in Natural Language Processing (Grant EP/S022481/1), School of Informatics, and School of Philosophy, Psychology & Language Sciences, the University of Edinburgh. We would like to thank Dr. Hao Tang and Dr. Tianzi Wang for useful discussion.
Appendix A ABX Conditions
ABX type on $(A=X\neq B)$ by $(A=B=X)$ across $(A\neq X)$
Accent accent word speaker
Speaker speaker word, accent
Phone phone phone, context speaker
Table 2: Different ABX conditions for selecting $(a,b,x)$ triplets to evaluate accent, speaker, and phonetic information accessibility.
: context refers to the previous and next phones.
[…]
Appendix B Selected ABX Accent-word Combination
After obtained a list of (accent $A$ and $B$, word) combinations with scores measured on GenAID features, we take the first p% of combination with lowest scores, which captures the most typical accent-specific pronunciation differences. A grid search on $p\%$ is performed over 2.5%, 5%, 10%, 15% and 20%, and we reported results at 10%, which is the lowest percentage where the score starts to show a converging trend.
We take the first 25 selected combination as an example, presented in Table 3.
Accent Word Accentb
SouthernEnglish first Scottish
NorthernEnglish first Scottish
SouthernEnglish Scottish Irish
SouthernEnglish however Scottish
SouthernEnglish work Scottish
NorthernEnglish however Scottish
Irish last Scottish
Scottish Scottish Irish
NorthernEnglish Scottish Irish
SouthernEnglish work Irish
NorthernEnglish their Scottish
SouthernEnglish first Irish
NorthernEnglish however Irish
SouthernEnglish however Irish
SouthernEnglish their Scottish
SouthernEnglish never Scottish
SouthernEnglish year Scottish
SouthernEnglish next Irish
SouthernEnglish Mr. Scottish
NorthernEnglish year Scottish
NorthernEnglish other Irish
SouthernEnglish other Irish
Scottish first SouthernEnglish
SouthernEnglish from Irish
Scottish first NorthernEnglish
Table 3: An example of the 25 triplet of (Accent AA, Accent $B$, word) with lowest ABX error rates, selected based on GenAID Zhong et al. (2025b) features on VCTK, applying to evaluate DSRTs accent ABX error rate in the test.
The table reveals a few well-established phonetic dimensions that underlie accent discrimination, successfully captured by our data-driven method. First, rhoticity emerges as a dominant cue. Words such as first and work sharply distinguish Southern or Northern English from Scottish and Irish accents, reflecting the systematic presence of post-vocalic /r/ in Scottish and Irish English and its absence in Southern English. The table shows that our metric takes good advantage of using rhoticity difference.
Second, differences in vowel quality, particularly in stressed lexical vowels, play a central role. For instance, first and year involve the near-front vowel spaces, which are realized as long centralized vowels in Southern English but shift toward more open or fronted realizations in Scottish and Irish English.
Third, consonantal realization, especially of /t/, contributes substantially to discrimination. Words such as scottish and last contain environments where Scottish English strongly favors glottal or lenited realizations of /t/, in contrast to the more alveolar realizations found in Southern and Irish English.
Finally, patterns of vowel reduction and weak-form realization further differentiate accents. Function words such as however, other, and from show strong schwa reduction in Southern English, whereas Irish and Scottish English tend to preserve fuller vowel qualities. These differences in reduction strategies, particularly in unstressed positions, amplify accent contrasts and explain why relevant accent-word combiantion are picked in the table.
From the analysis above, we can see the accent-word combination selected to evaluate accent ABX has a strong grounding in phonetic feature. The low accent accessibility results shown in Figure 4 and 3 authentically reflect the lack of accessible accent encoding in current speech representations and DSRTs.
Appendix C Accent and Speaker Partition in Cross-Accent Voice Conversion
See Table 4.
Accent Group Accent Region Split Speaker IDs
Seen during HiFiGAN training & Accent ABX word selection Southern English Train/Valid p225, p226, p228, p229, p231, p232, p239, p240, p243, p250, p254, p257, p258
Test p268, p273, p274, p276
Northern English Train/Valid p227, p230, p233, p236, p244, p256, p259, p267, p269, p270, p278, p279
Test p277, p282, p286, p287
Scottish Train/Valid p234, p237, p241, p246, p247, p249, p252, p255, p260, p262, p263, p271, p272, p275, p281
Test p264, p265, p284, p285
Irish Train/Valid p245, p266, p283, p288, p295
Test p298, p313, p340, p364
Likely seen during RepCodec training Canadian Test p316, p317, p343, p363
American Northeast Test p315, p339, p360, p361
American Midwest Test p311, p333, p334, p341
American South Test p301, p308, p310, p345
American West Test p294, p299, p300, p318
Unseen accents Northern Irish Test p292, p293, p304, p351
South African Test p314, p323, p336, p347
Indian Test p248, p251, p376
Oceanian Test p326, p335, p374
Table 4: Accent grouping and train/test splits for VCTK speakers. Accents are grouped by whether they were seen during HiFiGAN training & Accent ABX word selection, likely seen during RepCodec training, or unseen during training.
Appendix D Full Evaluation Results
Due to page limits, we report additional results here. Accent COS-SIM (to tar), Speaker COS-SIM (to tar), and WER exhibit reverse/similar trend, compared to Accent COS-SIM (to src), Speaker COS-SIM (to src), and PPG Distance, and are thus omitted in Figure 2 and 3. Information accessibility evaluation results across codebook sizes are also reported here in Figure 3.
Figure 4: Accent, speaker, and phonetic information in different DSRTs across layers and representations.
Full version of Figure 2.
Figure 5: Accent, speaker, and phonetic information in different DSRTs across codebook sizes.
Full version of Figure 3.

### v2

@@
-Rethinking Discrete Speech Representation Tokens for Accent Generation
+Zhong Wang Richmond Bell
 Abstract
@@
 We propose a unified evaluation framework that measures both accessibility of accent information via a novel Accent ABX task and recoverability via cross-accent Voice Conversion (VC) resynthesis.
-Using this framework, we analyse DSRTs derived from a variety of speech encoders.
-Our results reveal that accent information is substantially reduced when ASR supervision is used to fine-tune the encoder, but cannot be effectively disentangled from phonetic and speaker information through naive codebook size reduction.
-Based on these findings, we propose new content-only and content-accent DSRTs that significantly outperform existing designs in controllable accent generation.
-Our work highlights the importance of accent-aware evaluation and provides practical guidance for designing DSRTs for accent-controlled speech generation.
-Rethinking Discrete Speech Representation Tokens for Accent Generation
-Jinzuomu Zhong1thanks: Equal contribution., Yi Wang [^1] footnotemark: 1, Korin Richmond1, Peter Bell1, 1Centre for Speech Technology Research, University of Edinburgh Correspondence: {jinzuomu.zhong, korin.richmond}@ed.ac.uk
+Using this framework, we analyse DSRTs derived from several widely used speech representations.
+Our results reveal that: (1) choice of layers has the most significant impact on retaining accent information, (2) accent information is substantially reduced by ASR supervision; (3) naive codebook size reduction cannot effectively disentangle accent from phonetic and speaker information.
+keywords:
+accent, discrete speech representation, voice conversion, accent conversion
 1 Introduction
@@
 By compressing speech signals or learned continuous representations into a finite set of discrete units, discrete speech tokens enable efficient storage of complex acoustic information while remaining inherently compatible with LLM architectures.
-As a result, they are rapidly becoming established as a foundational representation across a wide range of speech tasks (Guo et al., 2025).
-Discrete speech representation tokens (DSRTs), quantised from speech representations, have consequently seen huge success in various speech generation tasks, including Speech Language Models (SpeechLMs) (Cui et al., 2025), Zero-Shot Text-to-Speech (ZS-TTS) (Kharitonov et al., 2023), Speech-to-Speech Translation (S2ST) (Lee et al., 2022) and full-duplex dialogue systems (Nguyen et al., 2023).
-Despite the significant progress brought by DSRTs, the representation of a speaker’s accent is largely overlooked in the design, evaluation, and application of the tokens.
-Prior user perception studies have demonstrated a similarity–attraction effect, whereby listeners prefer accents similar to their own in spoken interaction (Dahlbäck et al., 2007).
-Unfortunately, ZS-TTS systems are shown to hallucinate accents that differ from those of the reference speaker (Zhong et al., 2025b), while spoken dialogue systems still lack broad adoption of benchmarks or evaluation settings that consider appropriate, accent-controlled speech generation (Cheng et al., 2025).
-Although some recent zero-shot TTS (ZS-TTS) systems have demonstrated some ability to mimic or control the accent of generated speech using discrete speech representation tokens (DSRTs), how much accent information is encoded in these tokens remains unexplored (Du et al., 2024; Zhang et al., 2025; Wang et al., 2025).
-Existing claims – such as that naive codebook size adjustment (Zhang et al., 2025) or ASR supervision (Du et al., 2024) can facilitate controlled accent generation – lack systematic investigation.
+As a result, they are rapidly becoming established as a foundational representation across a wide range of speech tasks [guo2025recent].
+Directly compressing raw speech signals, however, tends to produce a representation that preserves abundant low-level acoustic detail, making phonetic content difficult to access.
+Such an approach works well for waveform reconstruction, but places a heavier modelling burden on downstream speech generation tasks [wang2025maskgct].
+To address this, discrete speech representation tokens (DSRTs), which are quantised from speech representations, are often used.
+DSRTs have seen huge success in various speech generation tasks, including Speech Language Models (SpeechLMs) [cui2025recent], Zero-Shot Text-to-Speech (ZS-TTS) [kharitonov2023speak], Speech-to-Speech Translation (S2ST) [lee2022direct] and full-duplex dialogue systems [nguyen2023dgslm].
+Despite the significant progress brought by DSRTs, the representation of a speaker's accent is largely overlooked in the design, evaluation, and application of the tokens.
+Prior user perception studies have demonstrated a similarity–attraction effect, whereby listeners prefer accents similar to their own in spoken interaction [dahlback2007similarity].
+Unfortunately, ZS-TTS systems are shown to hallucinate accents that differ from those of the reference speakers [zhong2025accentbox], while spoken dialogue systems still lack broad adoption of benchmarks or evaluation settings that consider appropriate, accent-controlled speech generation [cheng2025voxdialogue].
+Although some recent ZS-TTS systems have demonstrated some ability to mimic or control the accent of generated speech using DSRTs, how much accent information is encoded in these tokens remains unexplored [du2024cosyvoice, zhang2025vevo, wang2025maskgct].
+Existing claims – such as that naive codebook size adjustment [zhang2025vevo] or ASR supervision [du2024cosyvoice] can facilitate accent control and generation – lack systematic investigation.
+Crucially, the accent information actually preserved in these DSRTs remains unquantified, leaving it unclear whether the observed accent generation capabilities in ZS-TTS stem from the representations themselves or are merely a byproduct of large-scale pretraining.
 To this end, we ask the following research questions: (1) How do different design choices in DSRTs influence the amount of accent information they encode?
@@
 in voice conversion (VC)?
-Existing DSRT evaluation frameworks focus on phonetic and speaker information and rarely consider accent (Polyak et al., 2021).
-We extend the investigation of accessibility to accent information, with an accent ABX method we proposed that evaluates the discriminability of representations for words uttered in different accents.
-Additionally, we propose to evaluate the recoverability of accent, speaker, and phonetic information by cross-accent VC, resynthesising from DSRTs from source speakers and speaker IDs from target speakers of different accents.
-Using the proposed framework, we observe the following key findings: (1) Accent information is most prominent in mid-early layers of HuBERT, different from speaker or phonetic information distribution.
+To answer the research questions, we propose a framework for evaluating DSRTs from both the recoverability and accessibility perspectives, in terms of accent, speaker, and phonetic information encoded.
+Existing DSRT evaluation frameworks focus on phonetic and speaker information and rarely consider accent [polyak2021speech].
+To the best of our knowledge, this is the first work to incorporate accent in evaluating DSRTs, and to apply ABX to the study of accent.
+Specifically, we extend the existing investigation of accessibility to accent information, with a novel accent ABX method that evaluates the discriminability of representations for words uttered in different accents.
+Additionally, we propose to evaluate the recoverability of accent, speaker, and phonetic information by cross-accent VC, resynthesising with DSRTs from source speakers and speaker IDs from target speakers of different accents as input.
+Using this proposed framework, we observe the following key findings: (1) Accent information is most prominent in mid-early layers of HuBERT, different from speaker or phonetic information distribution.
 (2) Naive adjustment of codebook sizes provides only very limited disentanglement of accent, speaker and content information.
-(3) Predominant design of DSRTs for speech generation (quantising a later layer in a speech representation model, or using ASR supervison) discards most accent information.
-Our main contributions are:
-  • To the best of our knowledge, this is the first work to incorporate accent in evaluating DSRTs, and to apply ABX to a study of accent.
-  • Based on our findings, we propose quantisation schemes that are appropriate for accent-preserving VC – preserving source speaker accent – and accent-adaptive VC – adapting to target speaker accent – which achieve superior performance to existing approaches.
-  • Evaluation pipeline and proposed DSRTs will be open-sourced in the future.
-Figure 1: Proposed pipeline for evaluating the recoverability and accessibility of accent, speaker, and phonetic information in various Discrete Speech Representation Tokens (DSRTs).
+(3) Predominant design of DSRTs for speech generation (quantising a later layer in a speech representation model, or using ASR supervision) discards most accent information.
+These findings challenge existing claims using naive codebook size adjustment [zhang2025vevo] or ASR supervision [du2024cosyvoice] for accent control and generation.
+Finally, based on our findings, we propose DSRT design choices that are appropriate for both accent-preserving VC – preserving source speaker accent, and accent-adaptive VC – adapting to target speaker accent.
+Both objective and subjective evaluation show superior performance to existing approaches.
+We will release links to code and demo in the final version.
 2 Related Work
 2.1 Discrete Speech Tokens
-Discrete speech tokens can be largely categorised into two groups: semantic tokens [^1] A severe misnomer as these tokens contain primarily phonetic information and little semantic information (Wells et al., 2022; Choi et al., 2024).
+Discrete speech tokens can be largely categorised into two groups: semantic tokens [^1] A severe misnomer as these tokens contain primarily phonetic information and little semantic information [wells2022phonetic, choi2024self].
 and acoustic tokens.
 We refer to semantic tokens as Discrete Speech Representation Tokens (DSRTs) here, naming them strictly by how these tokens are designed and trained, rather than what information some researchers assume them to possess.
-DSRTs are tokens obtained by quantising learned speech representations, such as HuBERT (Hsu et al., 2021) and w2v-BERT (Chung et al., 2021); while acoustic tokens are obtained by quantising waveforms.
-Common quantisation techniques include k-means, Vector Quantisation (VQ) (van den Oord et al., 2017), Residual Vector Quantisation (RVQ) (Zeghidour et al., 2021), and Finite-State Quantisation (FSQ) (Mentzer et al., 2024), with k-means, VQ, and FSQ primarily used for DSRTs, VQ and RVQ primarily used for acoustic tokens.
+DSRTs are tokens obtained by quantising learned speech representations, such as HuBERT [hsu2021hubert] and w2v-BERT [chung2021w2vbert]; while acoustic tokens are obtained by directly quantising waveforms.
+Common quantisation techniques include k-means, Vector Quantisation (VQ) [vandenoord2017neural], Residual Vector Quantisation (RVQ) [zeghidour2021soundstream], and Finite-State Quantisation (FSQ) [mentzer2024finite], with k-means, VQ, and FSQ primarily used for DSRTs, VQ and RVQ primarily used for acoustic tokens.
 2.2 Investigation of Speech Representations
-In terms of of continuous speech representations, most previous work has focused on accessibility, finding that phonetic information peaks in intermediate layers, while acoustic information dominates earlier layers (Choi et al., 2024; Pasad et al., 2021).
+In terms of of continuous speech representations, most previous work has focused on accessibility, finding that phonetic information peaks in intermediate layers, while acoustic information dominates earlier layers [choi2024self, pasad2021layer].
 After finetuning on the ASR task, it has been observed deeper layers become dominated by task-specific information, exhibiting higher phonetic discrimination.
 In the investigation of DSRTs, Wells et al.
-(2022) find that DSRTs obtained by k-means on hubert-base-ls960 correspond to sub-phonetic events.
-Yeh and Tang (2024), using their information completeness and accessibility framework, found that DSRTs obtained from hubert-base-ls960 layer 4 contain more accessible pitch and speaker information, but less accessible phone information, than layer 9.
+[wells2022phonetic] find that DSRTs obtained by k-means on hubert-base-ls960 correspond to sub-phonetic events.
+Yeh et al.
+[yeh2024estimating], using their information completeness and accessibility framework, found that DSRTs obtained from hubert-base-ls960 layer 4 contain more accessible pitch and speaker information, but less accessible phone information, than layer 9.
 Phonetic and speaker information in both layers consistently decrease with lower bitrates of RVQ codes, refuting claims that content and speaker disentanglement can be achieved by VQ.
 2.3 DSRTs in ZS-TTS and SpeechLMs
-In ZS-TTS, researchers found that DSRTs provide more accessible phonetic information than acoustic tokens do, while suffering information loss such as timbre and acoustic environment (Wang et al., 2025).
-Motivated by such findings, many state-of-the-art (SOTA) models utilise a hierarchical approach, first predicting DSRTs from text or source speech and then predicting acoustic tokens from DSRTs (Du et al., 2024; Lee et al., 2025; Wang et al., 2025; Zhang et al., 2025).
-In SpeechLMs, researchers seek to combine DSRTs with acoustic tokens for both accessible and complete/recoverable information, such as SpeechTokenizer (Zhang et al., 2024) and Mimi tokeniser (Défossez et al., 2024).
-When designing such DSRTs, ZS-TTS researchers have claimed that are unverified or even contradicted by speech representation investigations.
-Zhang et al.
-(2025) choose layer 18 of hubert-large-ll60k and claim that reducing the codebook size in VQ leads to natural disentanglement of speaker, style (accent and emotion by their definition) and content, with a codebook size of 32 containing only content, and a codebook size of 8,192 containing only content and style.
-This claim not only contradicts the findings of Yeh and Tang (2024) that VQ or k-means with 1024 codebook size still retain significant speaker information, but also lacks verification in terms of accent and emotion.
-Du et al.
-(2024) propose to use “supervised semantic token”, obtained by injecting FSQ in an internal ASR model encoder; how ASR pretraining affects the information in discrete tokens remains unclear.
+In ZS-TTS, researchers found that DSRTs provide more accessible phonetic information than acoustic tokens do, while suffering information loss such as timbre and acoustic environment [wang2025maskgct].
+Motivated by such findings, many state-of-the-art (SOTA) models utilise a hierarchical approach, first predicting DSRTs from text or source speech and then predicting acoustic tokens from DSRTs [du2024cosyvoice, lee2025hierspeechpp, wang2025maskgct, zhang2025vevo].
+In SpeechLMs, researchers seek to combine DSRTs with acoustic tokens for both accessible and complete/recoverable information, such as SpeechTokenizer [zhang2024speechtokenizer] and Mimi tokeniser [defossez2024moshi].
+When designing such DSRTs, ZS-TTS researchers have introduced some claims that are either unverified or contradicted by speech representation investigations.
+Vevo [zhang2025vevo] chooses layer 18 of hubert-large-ll60k and claims that reducing the codebook size in VQ leads to natural disentanglement of speaker, style (accent and emotion by their definition) and content, with a codebook size of 32 containing only content, and a codebook size of 8,192 containing only content and style.
+This claim not only contradicts the findings in [yeh2024estimating] that VQ or k-means with 1024 codebook size still retain significant speaker information, but also lacks verification in terms of accent and emotion.
+CosyVoice [du2024cosyvoice] proposes to use ``supervised semantic token'', obtained by injecting FSQ in an internal ASR model encoder, for content consistency in the generated speech; how ASR pretraining affects accent information in discrete tokens remains unclear.
 2.4 Accent Control and Generation
-Unfortunately, ZS-TTS speech generation systems suffer accent hallucination, whereby they generate a hallucinated synthetic accent that deviates from the reference or prompt speech (Zhong et al., 2025b).
-Recent systems that have been specially built for accent generation have not yet utilised DSRTs along with powerful LLM architectures (Zhong et al., 2025b; Xinyuan et al., 2025).
+Unfortunately, ZS-TTS speech generation systems suffer accent hallucination, whereby they generate a hallucinated synthetic accent that deviates from the reference or prompt speech [zhong2025accentbox].
+Recent systems that have been specially built for accent generation have not yet utilised DSRTs along with powerful LLM architectures [zhong2025accentbox, xinyuan2025scalable].
 General ZS-TTS systems are reported to have certain accent control capabilities, but there has been little or no investigation about accent information in DSRTs.
@@
 2.5 ABX Evaluation
-[…]
-Previous work introduced Minimal Pair ABX (MP-ABX) (Schatz et al., 2013), where instances differ only in a central phonetic unit while sharing the same phonetic context.
-Prior studies (Borsos et al., 2023; Wang et al., 2025; Zhang et al., 2025) have used this metric to assess the phonetic discriminability of DSRT tokens.
-In addition, ABX has been extended to probe speaker-related information by defining triplets that contrast speaker identity under controlled linguistic content (de Seyssel et al., 2022).
+The ABX error rate is a distance-based, model-agnostic evaluation metric, assessing whether the representations pull data of the same category (e.g.
+sample $x=$ [bIt], another sample $a=$ [bIt], of the same triphone category) closer and push dissimilar data (e.g.
+$x=$ [bIt], $b=$ [bi:t]) farther in representational space.
+Previous work introduced Minimal Pair ABX (MP-ABX) [schatz2013evaluating], where instances differ only in a central phonetic unit while sharing the same phonetic context.
+Prior studies [wang2025maskgct, borsos2023audiolm, zhang2025vevo] have used this metric to assess the phonetic discriminability of DSRT tokens.
+In addition, ABX has been extended to probe speaker-related information by defining triplets that contrast speaker identity under controlled linguistic content [de2022language].
 Nevertheless, to the best of our knowledge, no prior work has systematically employed machine ABX to reveal accent discriminability in speech representations or DSRTs.
 3 Method
-Inspired by the perspective of separately measuring information completeness and accessibility Yeh and Tang (2024), we focus on evaluating DSRTs from both aspects, but in the context of accent generation.
+Figure 1: Proposed pipeline for evaluating the recoverability and accessibility of accent, speaker, and phonetic information in various Discrete Speech Representation Tokens (DSRTs).
+Inspired by the perspective of separately measuring information completeness and accessibility [yeh2024estimating], we focus on evaluating DSRTs from both aspects, but in the context of accent generation.
 Since completeness cannot be measured by resynthesis, we introduce recoverability as a task-grounded measure of how much accent, speaker, and phonetic information encoded in DSRTs can be recovered in resynthesised speech.
@@
 Figure 1 provides an overview of the proposed pipeline.
-First, we obtain DSRTs from several commonly used speech representations using RepCodec Huang et al.
-(2024) with Vector Quantisation (VQ) van den Oord et al.
-(2017) as the quantiser (see Section 3.1).
-For each DSRT configuration, we then train a unit-to-speech resynthesis model using HiFiGAN Polyak et al.
-(2021).
+First, we obtain DSRTs from several commonly used speech representations using RepCodec [huang2024repcodec] with Vector Quantisation (VQ) [vandenoord2017neural] as the quantiser (see Section 3.1).
+For each DSRT configuration, we then train a unit-to-speech resynthesis model using HiFiGAN [polyak2021speech].
 To assess information recoverability, we infer unit-to-speech models with DSRTs from a source speaker and speaker ID from a target speaker with a different accent, thereby conducting cross-accent VC.
@@
 We choose three speech representation models to obtain Discrete Speech Representation Tokens (DSRTs): HuBERT; HuBERT finetuned for ASR (HuBERT-ft); and Whisper.
-We include HuBERT Hsu et al.
-(2021) as it provides the most widely used speech representations in various speech tasks such as SpeechLM and TTS Guo et al.
-(2025); Cui et al.
-(2025).
+We include HuBERT [hsu2021hubert] as it provides the most widely used speech representations in various speech tasks such as SpeechLM and TTS [guo2025recent, cui2025recent].
 Recently, there has been a growing trend toward using ASR-based speech representations for obtaining DSRTs.
-Therefore, we include HuBERT-ft and Whisper as two representative ASR-based representations, with Encoder-only and Encoder-Decoder architectures, respectively.
-Following RepCodec Huang et al.
-(2024), we use VQ-VAE van den Oord et al.
-(2017) to discretise the speech representations.
+Therefore, we include HuBERT-ft and Whisper [radford2023robust] as two representative ASR-based representations, with Encoder-only and Encoder-Decoder architectures, respectively.
+All three speech representations investigated, despite being at least three years old, are still widely used in current speech LLMs, evidenced by their usage in speech tokenisers of 19 systems [cui2025recent, Table 1].
+Following RepCodec [huang2024repcodec], we use VQ-VAE [vandenoord2017neural] to discretise the speech representations.
 The model consists of three modules: Encoder and Decoder, which are both convolution layers with residual paths, and a Vector Quantisation module that quantises latent representations from Encoder output into a series of discrete tokens.
 The model is trained to reconstruct the speech representation, with additional Exponential Moving Average (EMA) optimisation to gradually update the codebook.
-For model details, we refer readers to the original paper Huang et al.
-(2024).
+For model details, we refer readers to the original paper [huang2024repcodec].
 3.2 Cross-Accent Voice Conversion
-After extracting DSRTs, we train a unit-to-speech HiFiGAN model for each DSRT configuration, following Polyak et al.
-(2021).
+After extracting DSRTs, we train a unit-to-speech HiFiGAN model for each DSRT configuration, following [polyak2021speech].
 In prior work, information recoverability is typically evaluated using resynthesis or Voice Conversion (VC) on General American English datasets, with evaluation focusing primarily on intelligibility and speaker similarity, while accent is largely ignored.
@@
 We then perform cross-accent VC in inference, by conditioning the model on DSRTs from a source speaker and a target speaker ID with a different accent.
-Afterwards, during evaluation, following Zhong et al.
-(2025a), we explicitly assess accent, speaker, and phonetic similarity in the converted speech.
+Afterwards, during evaluation, following [zhong2025pairwise], we explicitly assess accent, speaker, and phonetic similarity in the converted speech, against ground truth reference clips.
 Utterance-level embeddings from Accent Identification (AID) and Speaker Verification (SV) models are used to calculate cosine similarity, which serves as a proxy for perceived accent and speaker similarity.
@@
 We choose ABX as a model-free method to estimate the accessibility of accent information, as well as phone and speaker information in DSRTs.
-In accent ABX, instances $a$ and $x$ share the same accent, while instance $b$ differs in accent.
-Triplets $(a,b,x)$ are constructed to share identical lexical content rather than phonetic context, as accent distinctions arise from accent-dependent realizations of the same words or lexical units, which may involve different phone sequences.
-We further require that $a$, $b$, and $x$ come from different speakers to avoid trivial similarity between $a$ and $x$ due to shared speaker identity, e.g.: a triplet $(a,b,x)$ consists of: $x=$ “water” spoken by a Scottish speaker1, $a=$ “water” by a Scottish speaker2, and $b=$ “water” by a Southern English speaker3.
-We also adopt speaker ABX and phone ABX as contrastive evaluations. For detailed triplet selection criteria, see Appendix A.
-In practice, words vary in both accent-induced pronunciation variation and speaker-specific pronunciation patterns.
-While standard ABX aggregates scores over all valid triplets, we adopt a word selection scheme to improve the sensitivity of accent ABX.
-We first select the 100 most frequent words in the training data, compute ABX scores for each (accent $A$ and $B$, word) combination, and retain the 10% with the lowest scores as the most accent-discriminative.
-The final accent ABX score is obtained by averaging their ABX scores on the test set.
-The combination selection is performed using continuous features from GenAID Zhong et al.
-(2025b), a supervised accent identification model fine-tuned from XLSR Babu et al.
-(2021), to ensure effective word selection and a fair comparison among DSRTs.
-See Appendix B for details.
-(a) Cross-accent VC evaluation results for information recoverability.
-(b) ABX evaluation results for information accessibility.
-Figure 2: Accent, speaker, and phonetic information in different DSRTs across layers and representations.
-All codebook sizes and code dimensions are set to 1024 for fair comparison.
-(See Figure 4 in Appendix D for additional results from cross-accent VC.)
-4 Experiments
-4.1 Obtaining DSRTs
-We extract continuous speech representations from various layers of HuBERT222hubert-large-ll60k, available here, HuBERT-ft333hubert-large-ls960-ft, available here and Whisper444whisper-meidum.en, available here.
-All three speech representation models share the same Encoder architecture (24 Transformer layers) and are trained on English-only data.
-Here, we focus on investigating how layer choice and ASR pretraining affect the information encoded in speech representations, and leave the investigation of how multilingual pretraining or larger-scale pretraining data/model to future work.
-The extracted representations are then discretised using RepCodec [^5] https://github.com/mct10/RepCodec, which we train on the train-clean-100 subset of LibriSpeech [^6] https://www.openslr.org/12 following similar practice in Huang et al.
-(2024) and Zhang et al.
-(2025), with various codebook sizes.
-This subset contains 100 hours of speech with "accents closer to US English" Panayotov et al.
-(2015).
-RepCodec is trained for 200,000 steps with the same hyperparameters in Huang et al.
-(2024).
-4.2 Evaluation Dataset
-To evaluate information recoverability and accessibility with respect to accent, we use the VCTK corpus [^7] https://datashare.ed.ac.uk/handle/10283/3443, which provides relatively broad coverage of native English accents.
-Based on the provided country labels and detailed region descriptions, we manually group speakers into 13 accent regions, as some coarse country-level labels (e.g.
-“US”, “England”) conflate multiple distinct accent varieties.
-We withhold 3–4 speakers from each accent region (50 speakers in total, with balanced gender and accent distribution) for testing.
-For the remaining speakers, we use 4 accent regions that have sufficient speakers (45 speakers in total) for training unit-to-speech HiFiGAN in Section 3.2 and for selecting accent-discriminative words in Section 3.3.
-Out of the 13 accent regions, 4 are seen in HiFiGAN training and accent ABX word selection, 5 are likely seen by RepCodec in LibriSpeech, and the remaining 4 are unseen at any stage, allowing for analysis on the generalisability of DSRTs across accents.
-See Appendix C for details of accent grouping and speaker splits.
-Figure 3: Accent, speaker, and phonetic information recoverability in different DSRTs across codebook sizes.
-(See Figure 5 in Appendix D for additional results from cross-accent VC and ABX.)
-4.3 Evaluating Information Recoverability
-To evaluate information recoverability, we train unit-to-speech HiFiGAN models [^8] https://github.com/facebookresearch/speech-resynthesis for 100,000 steps with the same hyperparameters as Polyak et al.
-(2021), using 4 accent regions (45 speakers), as described in Section 4.2.
-During inference, we perform cross-accent VC using source (src) DSRTs from 12 accent regions (46 speakers, excluding Southern English), together with speaker IDs from 4 Southern English speakers as target (tgt) voices.
-We fix the target voices to one accent region to control for target-side accent variability, enabling a clearer analysis of how accent information is preserved in the src DSRTs.
-For objective evaluation of the converted speech, we adopt metrics recommended by Zhong et al.
-(2025a).
-(1) Accent similarity: We calculate cosine similarity of accent embeddings (Accent COS-SIM) extracted from GenAID [^9] https://github.com/jzmzhong/GenAID Zhong et al.
-(2025b).
-GenAID is trained with explicit speaker-accent disentanglement, preserving little speaker information in the extracted accent embeddings.
-(2) Speaker similarity: We calculate cosine similarity of speaker embeddings (Speaker COS-SIM) extracted from WavLM101010wavlm-base-plus-sv, available here Chen et al.
-(2022).
-These speaker embeddings encode some accent information, evidenced by improved performance in AID when transfer learned from SV models Zuluaga-Gomez et al.
-(2023).
-(3) Phonetic similarity: We calculate Jensen-Shannon distance between aligned PPGs [^11] https://github.com/interactiveaudiolab/ppgs (PPG Distance), extracted from a phone recognition model Churchwell et al.
-(2024).
-It is worth noting that PPG distance, primarily designed for pronunciation distance, is sensitive to accent information, as accent can be characterised partially as different realisations of the same phoneme.
-(4) Intelligibility: We calculate Word Error Rate (WER) using whisper-medium.en transcriptions, compared against ground-truth transcriptions.
-Despite its evident accent bias Jahan et al.
-(2025), WER is included here due to its broad adoption in previous DSRT evaluation frameworks to reflect how DSRTs are commonly selected.
-For Accent COS-SIM, Speaker COS-SIM, and PPG distance, we use the first 24 utterances per speaker, which is an elicitation paragraph shared across speakers.
-In this way, we allow for similarity/distance calculation to either source speech that provides DSRTs or target speech that provides the speaker ID of the same content, removing the influence of content on these metrics.
-For WER, we randomly choose 24 utterances per speaker to cover diverse contents.
-For visualisation, we annotate each metric with a practical lower bound (chance-level performance) and a practical upper bound (best possible performance), which anchor the visual scale and facilitate interpretation of values.
-For Accent COS-SIM, PPG, and WER, the lower bound measures the source speaker ground truth (GT) with respect to target speaker GT, and the upper bound measures the source speaker copy-synthesis with respect to source speaker GT.
-For Speaker COS-SIM, the lower bound measures similarity between source speaker GT and target speaker GT, while the upper bound measures similarity between target speaker copy-synthesis and target speaker GT.
-4.4 Evaluating Information Accessibility
-Information accessibility is evaluated only on the accents seen during HiFiGAN training, ensuring sufficient numbers of speakers for each accent.
-As described in Section 3.3, a training set is required to select high-frequency words and accent-discriminative combinations, while a separate test set is used for ABX evaluation to prevent speaker leakage.
-Accents outside the four seen accents contain no more than four speakers each, severely limiting the number of valid ABX triplets.
-We follow the same training–test partition as in HiFiGAN experiments for simplicity.
-Accent ABX is conducted using this partition, whereas phone and speaker ABX are computed only on the test split, as they do not require a training phase.
-Test experiments used all utterances in the split, except for the first 24 utterances to avoid identical sentences across all train and test speakers, which could bias the ABX scores toward lexical-context matching instead of reflecting true accent or speaker discrimination.
-Phone-level alignment (phone boundaries and labels) required by phone ABX, were obtained through the Montreal Forced Aligner with the English Arpabet dictionary McAuliffe and Sonderegger (2024); Gorman et al.
-(2011).
-We treat the accent ABX score computed on HuBERT layer 6 continuous features as an upper bound, as this layer exhibits the strongest accent-related information.
-Similarly, we use the phonetic and speaker ABX scores computed on HuBERT-ft layers 3 and 9 continuous features as upper bounds, respectively, since these layers demonstrate the highest phonetic and speaker recoverability in Section 5.
-In addition, we evaluate the accent ABX error rate of the accent-discriminative GenAID features, which achieves a low error rate of 9.87%, validating the effectiveness of the proposed metric.
-Token type
-Representation
-& Layer
-Codebook
-size
-A-SIM
-A-SIM
-S-SIM
-[…]
-S-SIM
-[…]
-[…]
-Vevo Content HuBERT L18 32 0.8143 0.9369 0.6817 0.9651 0.1718 0.0824
-Content-style HuBERT L18 8192 0.8923 0.8888 0.6989 0.9499 0.1427 0.0425
-Proposed Content HuBERT-ft L18 256 0.8081 0.9509 0.6782 0.9667 0.1488 0.0523
-Content-accent HuBERT L9 8192 0.9541 0.8227 0.7529 0.9104 0.1265 0.0344
-Table 1: Results of proposed content and content-accent tokens, compared with content and content-style tokens from Vevo (Zhang et al., 2025).
-Differences in DSRT design choices are marked in red, with best performances marked in bold.
-*: For content tokens, we want little accent information recovered from DSRTs, i.e.
-low A-SIM (src) and high A-SIM (tar).
-For content-accent or content-style tokens, we want the opposite.
-5 Results
-5.1 Layer Choice Matters for Accent
-Layer choice has a prominent impact on the recoverability and accessibility of information in HuBERT DSRTs, shown in Figure 2 (dark green curve).
-Notably, accent, speaker and phonetic information are distributed differently across layers.
-(1) Accent information recoverability, reflected by Accent COS-SIM w.r.t.
-source speaker, is most prominent in mid–early HuBERT layers (L6 & L9), and decreases in earlier or latter layers, indicating that accent cues emerge after low-level acoustic processing but are progressively abstracted away in higher layers.
-(2) Speaker information is most recoverable in early HuBERT layers (L3), indicated by Speaker COS-SIM to the source speaker, and decreases monotonically with depth.
-Overall, speaker information is limited, as cross-accent VC aims to generate voice similar to that of the target speaker with speaker ID conditioning.
-(3) Phonetic information, measured by PPG Distance to the source speech, is most complete in middle layers (L9 & L12).
-Both earlier and later layers show reduced phonetic completeness, reflecting a trade-off between raw acoustic detail and higher-level abstraction.
-(4) Information accessibility, measured by ABX error rates (see Figure 2(b)), reveals weak accent accessibility for HuBERT across layers, despite clear inter-layer variations.
-Accent accessibility peaks at L12, which differs from the recoverability peak at L6 or L9.
-This discrepancy highlights the necessity of resynthesis-based recoverability evaluation when assessing DSRTs for accent generation: while ABX accessibility captures relative trends, it cannot reliably identify optimal layers for synthesis.
-5.2 ASR Supervision Removes Accent Information
-ASR supervision reduces the recoverability and accessibility of accent information in DSRTs across layers, shown in Figure 2 (blue curve for HuBERT-ft, orange curve for Whisper).
-(1) HuBERT consistently shows higher accent recoverability and accessibility than HuBERT-ft or Whisper, particularly in latter layers, which are more affected by ASR supervision.
-(2) Whisper exhibits a peak in accent and phonetic information at intermediate layers (L12 & L15), different from the distribution of information in HuBERT or HuBERT-ft.
-5.3 Reducing Codebook Size Provides Limited Disentanglement
-We then vary codebook sizes for three selected layers: HuBERT L18 (a common choice, as in Vevo Zhang et al.
-(2025)), HuBERT-ft L18 (a layer with little accent information while retaining high phonetic information), and HuBERT L9 (a layer with most accent information), shown in Figure 3.
-(1) Focusing on accent recoverability, varying codebook sizes has a markedly smaller effect than changing representation models or layers.
-For each layer, increasing codebook sizes from 32 to 2048 consistently improves accent and phonetic recoverability across all inspected DSRTs.
-However, the recoverability gap induced by codebook sizes variation within a layer is substantially smaller than that induced by changing layers at a fixed codebook sizes.
-(2) Speaker and phonetic recoverability also increase with larger codebook sizes.
-Taking HuBERT DSRTs as an example, Accent COS-SIM drops sharply as codebook sizes decreases from 1024 to 32, while PPG distance also rises 20.4%.
-Despite speaker identity being partially suppressed by the injected speaker ID during resynthesis, speaker recoverability also declines.
-The parallel degradation across accent, speaker, and phonetic metrics suggests minimal disentanglement when reducing codebook size.
-5.4 Limitations of Commonly Used DSRTs in Speech Generation
-Based on the above findings, we highlight the limitations of commonly used DSRTs here.
-(1) Vevo’s choice of using DSTRs based on HuBERT L18 with codebook size 32 as content tokens and codebook size 8,192 as content-style tokens Zhang et al.
-(2025) is suboptimal.
-Significant accent information is likely already lost and unrecoverable in HuBERT L18 - leading to insufficient accent information encoded in content-style tokens.
-Since naive reduction of codebook size cannot remove accent without harming phonetic information, content tokens achieve little accent information at the cost of intelligibility in resynthesised speech.
-(2) ASR-supervised DSRTs, adopted by Du et al.
-(2024), exhibit lower peak accent and phonetic recoverability than HuBERT DSRTs.
-Similar to content-style tokens, ASR-supervised DSRTs cannot encode accent sufficiently.
-5.5 Proposed Content and Content-Accent Tokens
-For better accent generation and control, we propose content-accent and content tokens for accent-preserving and accent-adaptive VC, respectively.
-A comparison of proposed tokens and Vevo tokens is presented in Table 1.
-Our proposed content/content-accent tokens achieve better performance than Vevo’s content/content-style tokens across all metrics, enabling better control of accent in speech generation.
-Demos of the converted speech will be available at [Link redacted].
-6 Discussion
-We observed a systematic delay of the optimal accessible layer relative to the recoverability peak.
-As discussed in (Yeh and Tang, 2024), earlier layers retain more complete acoustic information that can be effectively extracted by a strong HiFiGAN vocoder, whereas later layers gain phonetic discriminability from ASR-related objectives,
-Across both metrics, accent information follows a distinct yet intermediate trend between phonetic and speaker information.
-This aligns with linguistic intuition: accent is partially realised through systematic phonetic variation and partially through speaker-specific pronunciation traits.
-Layer selection enables accent control precisely because accent, speaker, and phonetic information are distributed differently across layers.
-Consequently, a single "style layer", as suggested by the content–style token design in Vevo, is insufficient to jointly capture multiple speech attributes such as content, accent, and emotion.
-A more thorough investigation of more speech attributes, and how their information are represented across multiple layers and representations, is therefore necessary to support better controllable speech generation.
-Finally, our findings provide potential explanations for the hallucinated accents by ZS-TTS systems Zhong et al.
-(2025b).
-Many ZS-TTS systems rely on representations (often deep or supervised layers) where accent recoverability is already attenuated, encouraging the model to guess accent patterns in inference or default to mainstream accents in training data.
-7 Conclusion
-In this work, we propose a framework for evaluating discrete speech representation tokens (DSRTs) from both recoverability and accessibility perspectives, in terms of accent, speaker, and phonetic information encoded, with a focus on providing guidance for accent generation and control.
-Based on the results using this framework on HuBERT- and Whipser-based DSRTs, we propose content and content-accent tokens, achieving more controllable accent in voice conversion (VC) than existing works.
-Limitations
-Although the framework we proposed guided us to develop a pair of DSRTs achieving good control over accent in VC, we acknowledge several limitations.
-First, we did not evaluate the proposed DSRTs in ZS-TTS settings, as training ZS-TTS systems requires substantial computational resources.
-Future work will extend our evaluation to ZS-TTS to assess the generalisability of our framework beyond voice conversion.
-Second, the current Accent ABX test shows low accent accessibility across all evaluated DSRTs.
-A supervised accent representation (GenAID Zhong et al.
-(2025b)) achieves much lower ABX error rate (9.87%), indicating the validity of the metric.
-More powerful probing models are needed to better reveal subtle accent information in SSL representations and DSRTs without explicit accent-discriminative training, so as to provide more precise guidance for accent applications.
-Better design of DSRTs is also needed to make accent information more accessible, and thus easier to learn by downstream speech generation models.
-Third, our experiments are limited to English SSL models and the VCTK dataset, which contains only L1 English speaker accents.
-We plan to extend the framework to multilingual representations and datasets involving L2 English accents and non-English languages to improve regional and language coverage.
-Finally, we focus on objective metrics to evaluate DSRTs and their effectiveness in VC.
-We haven’t included subjective evaluation and listening test, as Zhong et al.
-(2025a) have shown the effectiveness of the adopted objective metrics on evaluating accent generation, correlating with subjective listening tests results.
-Broader Impacts and Ethical Statements
-This work improves the transparency and controllability of accent information in discrete speech representation tokens (DSRTs).
-The proposed framework may help mitigate accent hallucination and unintended accent drift, benefiting speakers with underrepresented accents and potentially supporting applications such as speech therapy through accent- and voice-preserving speech generation.
-Our results further suggest that accent control can be partially achieved via informed representation and layer selection, potentially reducing reliance on additional large-scale model training.
-Although this work is intended for research use, enhanced accent controllability in speech generation could be misused for voice impersonation if deployed without safeguards.
-If extended to real-world or unseen speakers, practical systems should incorporate mechanisms to ensure speaker approval and to detect synthesized speech.
-We emphasise that our framework is designed to analyse and expose limitations of DSRTs rather than to enable unrestricted accent manipulation.
-This study exclusively uses existing publicly available datasets and pretrained models for research purposes, without collecting new data or releasing derived datasets.
-All data and pre-trained models are used under the following licenses: Creative Commons BY 4.0 License, Creative Commons BY-NC-ND 4.0 License, Apache License 2.0, and MIT License.
-Acknowledgments
-This work was supported in part by the UKRI AI Centre for Doctoral Training in Responsible and Trustworthy in-the-world Natural Language Processing (Grant EP/Y030656/1), UKRI Centre for Doctoral Training in Natural Language Processing (Grant EP/S022481/1), School of Informatics, and School of Philosophy, Psychology & Language Sciences, the University of Edinburgh.
-We would like to thank Dr.
-Hao Tang and Dr.
-Tianzi Wang for useful discussion.
-Appendix A ABX Conditions
+Table 1 lists the three different ABX conditions we use to evaluate accent, speaker, and phonetic information accessibility.
+Table 1: Different ABX conditions for selecting $(a,b,x)$ triplets to evaluate accent, speaker, and phonetic information accessibility. : context refers to the previous and next phones.
 ABX type on $(A=X\neq B)$ by $(A=B=X)$ across $(A\neq X)$
@@
 Speaker speaker word, accent
-Phone phone phone, context speaker
-Table 2: Different ABX conditions for selecting $(a,b,x)$ triplets to evaluate accent, speaker, and phonetic information accessibility.
-: context refers to the previous and next phones.
-[…]
-Appendix B Selected ABX Accent-word Combination
-After obtained a list of (accent $A$ and $B$, word) combinations with scores measured on GenAID features, we take the first p% of combination with lowest scores, which captures the most typical accent-specific pronunciation differences.
-A grid search on $p\%$ is performed over 2.5%, 5%, 10%, 15% and 20%, and we reported results at 10%, which is the lowest percentage where the score starts to show a converging trend.
-We take the first 25 selected combination as an example, presented in Table 3.
-Accent Word Accentb
+Phone phone context speaker
+When constructing $(a,b,x)$ triplets, a condition can be described as ABX ON the category of interest, BY the category of controlled constant among triplets, ACROSS the category of free variables among triplets.
+The classic Phone ABX select triplets ON phone BY context, where 1) $x$ and $a$ share the same phone, with $b$ of a different phone, and 2) $x$, $a$, and $b$ share identical phonetic context, e.g.
+bit [bIt] vs beat [bi:t].
+To evaluate how prominent phonetic information is, compared to other variables, such as speaker, most Phone ABX would also be conditioned ACROSS speaker, with $a$, $b$ coming from the same speaker, and $x$ from a different speaker, e.g.
+$x=$ [bIt] by speaker1, $a=$ [bIt] by speaker2, $b=$ [bi:t] by speaker2.
+In the proposed Accent ABX, instances $a$ and $x$ share the same accent, while instance $b$ differs in accent.
+Triplets $(a,b,x)$ are constructed to share identical lexical content rather than phonetic context, as accent distinctions arise from accent-dependent realisations of the same words or lexical units, which may involve different phone sequences.
+We further require that $a$, $b$, and $x$ come from different speakers to avoid trivial similarity between $a$ and $x$ due to shared speaker identity, e.g.
+a triplet $(a,b,x)$ comprosing $x=$ ``water'' spoken by a Scottish speaker1, $a=$ ``water'' by a Scottish speaker2, and $b=$ ``water'' by a Southern English speaker3.
+In practice, words uttered by different speakers differ due to both accent-induced pronunciation variation and speaker-specific pronunciation patterns.
+While standard ABX aggregates scores over all valid triplets, we adopt a word selection scheme to improve the sensitivity of Accent ABX.
+We first select the 100 most frequent words in the training data, compute ABX scores for each (accent $A$, accent $B$, word) combination, and retain the 10% with the lowest scores as the most accent-discriminative.
+The choice of most frequent words is to ensure there are enough triplets to sample from in Accent ABX, while the selection of most accent-discriminative combinations is to avoid conducting Accent ABX on words that are pronounced similarly in two distinct accents.
+The combination selection is performed using continuous features from GenAID [zhong2025accentbox], a supervised Accent Identification (AID) model finetuned from XLSR [babu2021xls], to ensure effective word selection and a fair comparison among DSRTs.
+Table 2: Samples of most accent-discriminative (Accent AA, Accent $B$, Word) combinations selected, for Accent ABX.
+Accent A Word Accent B
 SouthernEnglish first Scottish
@@
 Scottish first NorthernEnglish
-Table 3: An example of the 25 triplet of (Accent AA, Accent $B$, word) with lowest ABX error rates, selected based on GenAID Zhong et al.
-(2025b) features on VCTK, applying to evaluate DSRTs accent ABX error rate in the test.
-The table reveals a few well-established phonetic dimensions that underlie accent discrimination, successfully captured by our data-driven method.
+Table 2 shows the first 25 selected combinations as an example.
+The table reveals a few well-established phonetic dimensions that underlie accent differences, successfully captured by our data-driven selection scheme.
 First, rhoticity emerges as a dominant cue.
-Words such as first and work sharply distinguish Southern or Northern English from Scottish and Irish accents, reflecting the systematic presence of post-vocalic /r/ in Scottish and Irish English and its absence in Southern English.
+Words such as first and work sharply distinguish Southern or Northern English from Scottish and Irish accents, reflecting the systematic presence of post-vocalic [r] in Scottish and Irish English and its absence in Southern English.
 The table shows that our metric takes good advantage of using rhoticity difference.
 Second, differences in vowel quality, particularly in stressed lexical vowels, play a central role.
-For instance, first and year involve the near-front vowel spaces, which are realized as long centralized vowels in Southern English but shift toward more open or fronted realizations in Scottish and Irish English.
-Third, consonantal realization, especially of /t/, contributes substantially to discrimination.
-Words such as scottish and last contain environments where Scottish English strongly favors glottal or lenited realizations of /t/, in contrast to the more alveolar realizations found in Southern and Irish English.
-Finally, patterns of vowel reduction and weak-form realization further differentiate accents.
+For instance, first and year involve the near-front vowel spaces, which are realised as long centralised vowels in Southern English but shift toward more open or fronted realisations in Scottish and Irish English.
+Third, consonantal realisation, especially of [t], contributes substantially to discrimination.
+Words such as scottish and last contain environments where Scottish English strongly favours glottal or lenited realisations of [t], in contrast to the more alveolar realisations found in Southern and Irish English.
+Finally, patterns of vowel reduction and weak-form realisation further differentiate accents.
 Function words such as however, other, and from show strong schwa reduction in Southern English, whereas Irish and Scottish English tend to preserve fuller vowel qualities.
-These differences in reduction strategies, particularly in unstressed positions, amplify accent contrasts and explain why relevant accent-word combiantion are picked in the table.
-From the analysis above, we can see the accent-word combination selected to evaluate accent ABX has a strong grounding in phonetic feature.
-The low accent accessibility results shown in Figure 4 and 3 authentically reflect the lack of accessible accent encoding in current speech representations and DSRTs.
-Appendix C Accent and Speaker Partition in Cross-Accent Voice Conversion
-See Table 4.
-Accent Group Accent Region Split Speaker IDs
-Seen during HiFiGAN training & Accent ABX word selection Southern English Train/Valid p225, p226, p228, p229, p231, p232, p239, p240, p243, p250, p254, p257, p258
-Test p268, p273, p274, p276
-Northern English Train/Valid p227, p230, p233, p236, p244, p256, p259, p267, p269, p270, p278, p279
-Test p277, p282, p286, p287
-Scottish Train/Valid p234, p237, p241, p246, p247, p249, p252, p255, p260, p262, p263, p271, p272, p275, p281
-Test p264, p265, p284, p285
-Irish Train/Valid p245, p266, p283, p288, p295
-Test p298, p313, p340, p364
-Likely seen during RepCodec training Canadian Test p316, p317, p343, p363
-American Northeast Test p315, p339, p360, p361
-American Midwest Test p311, p333, p334, p341
-American South Test p301, p308, p310, p345
-American West Test p294, p299, p300, p318
-Unseen accents Northern Irish Test p292, p293, p304, p351
-South African Test p314, p323, p336, p347
-Indian Test p248, p251, p376
-Oceanian Test p326, p335, p374
-Table 4: Accent grouping and train/test splits for VCTK speakers.
-Accents are grouped by whether they were seen during HiFiGAN training & Accent ABX word selection, likely seen during RepCodec training, or unseen during training.
-Appendix D Full Evaluation Results
-Due to page limits, we report additional results here.
-Accent COS-SIM (to tar), Speaker COS-SIM (to tar), and WER exhibit reverse/similar trend, compared to Accent COS-SIM (to src), Speaker COS-SIM (to src), and PPG Distance, and are thus omitted in Figure 2 and 3.
-Information accessibility evaluation results across codebook sizes are also reported here in Figure 3.
-Figure 4: Accent, speaker, and phonetic information in different DSRTs across layers and representations.
-Full version of Figure 2.
+These differences in reduction strategies, particularly in unstressed positions, amplify accent contrasts and explain why relevant accent-word combinations are picked in the table.
+4 Experiments
+4.1 Obtaining DSRTs
+We extract continuous speech representations from various layers of HuBERT [^2] https://huggingface.co/facebook/hubert-large-ll60k, HuBERT-ft [^3] https://huggingface.co/facebook/hubert-large-ls960-ft and Whisper [^4] https://huggingface.co/openai/whisper-medium.en.
+All three speech representation models share the same Encoder architecture (24 Transformer layers) and are trained on English-only data.
+Here, we focus on investigating how layer choice and ASR pretraining affect the information encoded in speech representations, and leave the investigation of how multilingual pretraining or larger-scale pretraining data/model to future work.
+The extracted representations are then discretised using RepCodec [^5] https://github.com/mct10/RepCodec, which we train on the train-clean-100 subset of LibriSpeech [^6] https://www.openslr.org/12, similar to [zhang2025vevo, huang2024repcodec], with various codebook sizes.
+This subset contains 100 hours of speech with "accents closer to US English" [panayotov2015librispeech].
+RepCodec is trained for 200,000 steps with the same hyperparameters in [huang2024repcodec].
+4.2 Evaluation Dataset
+To evaluate information recoverability and accessibility with respect to accent, we use the VCTK corpus [^7] https://datashare.ed.ac.uk/handle/10283/3443, which provides relatively broad coverage of native English accents.
+Based on the provided country labels and detailed region descriptions, we manually group speakers into 13 accent regions, as some coarse country-level labels (e.g.
+``US'', ``England'') conflate multiple distinct regional accent varieties.
+We withhold 3–4 speakers from each accent region (50 speakers in total, with balanced gender and accent distribution) for testing.
+For the remaining speakers, we use 4 accent regions that have sufficient speakers (45 speakers in total) for training unit-to-speech HiFiGAN in Section 3.2 and for selecting accent-discriminative words in Section 3.3.
+Out of the 13 accent regions, 4 are seen in HiFiGAN training and Accent ABX word selection, 5 are likely seen by RepCodec in LibriSpeech given they are North American accents, and the remaining 4 are unseen at any stage, allowing for analysis on the generalisability of DSRTs across accents.
+4.3 Evaluating Information Recoverability
+To evaluate information recoverability, we train unit-to-speech HiFiGAN models [^8] https://github.com/facebookresearch/speech-resynthesis for 100,000 steps with the same hyperparameters as in [polyak2021speech], using 4 accent regions (45 speakers), as described in Section 4.2.
+During inference, we perform cross-accent VC using source (src) DSRTs from 12 accent regions (46 speakers, excluding Southern English), together with speaker IDs from 4 Southern English speakers as target (tgt) voices.
+We fix the target voices to one accent region to control for target-side accent variability, enabling a clearer analysis of how accent information is preserved in the source DSRTs.
+For objective evaluation of the converted speech, we adopt metrics recommended by [zhong2025pairwise].
+(1) Accent similarity: We calculate cosine similarity of accent embeddings (Accent COS-SIM or A-SIM) extracted from GenAID [^9] https://github.com/jzmzhong/GenAID [zhong2025accentbox].
+GenAID is trained with explicit speaker-accent disentanglement, preserving little speaker information in the extracted accent embeddings.
+(2) Speaker similarity: We calculate cosine similarity of speaker embeddings (Speaker COS-SIM or S-SIM) extracted from WavLM [^10] https://huggingface.co/microsoft/wavlm-base-plus-sv [chen2022wavlm].
+These speaker embeddings encode some accent information, evidenced by improved performance in AID when transfer learned from SV models [zuluagagomez2023commonaccent].
+(3) Phonetic similarity: We calculate Jensen-Shannon distance between aligned PPGs [^11] https://github.com/interactiveaudiolab/ppgs (PPG Distance), extracted from a phone recognition model [churchwell2024ppgs].
+It is worth noting that PPG distance, primarily designed for pronunciation distance, is sensitive to accent information, as accent can be characterised partially as different realisations of the same phoneme.
+(4) Intelligibility: We calculate Word Error Rate (WER) using whisper-medium.en transcriptions, compared against ground-truth transcriptions.
+Despite its evident accent bias [jahan2025unveiling], WER is included here due to its broad adoption in previous DSRT evaluation frameworks to reflect how DSRTs are commonly selected.
+We use the first 24 utterances per speaker, which is an elicitation paragraph shared across speakers, to calculate Accent COS-SIM, Speaker COS-SIM, and PPG distance.
+In this way, we allow for similarity/distance calculation to either source speech that provides DSRTs or target speech that provides the speaker ID of the same content, removing the influence of content on these metrics.
+As for WER, we randomly choose 24 utterances per speaker to cover diverse content.
+For clearer visualisation, we annotate each metric with a practical lower bound (chance-level performance) and a practical upper bound (best possible performance), which anchor the visual scale and facilitate interpretation of values.
+For Accent COS-SIM, PPG, and WER, the lower bound measures the source speaker ground truth with respect to target speaker ground truth, and the upper bound measures the source speaker copy-synthesis with respect to source speaker ground truth.
+For Speaker COS-SIM, the lower bound measures similarity between source speaker ground truth and target speaker ground truth, while the upper bound measures similarity between target speaker copy-synthesis and target speaker ground truth.
+Finally, we select several DSRTs to evaluate their performance in both accent-preserving VC and accent-adaptive VC, respectively, and conduct subjective listening tests on the converted speech.
+Three sets of 5-point Similarity MOS (SMOS) listening tests are conducted, with 15 listeners for each listening test, recruited on Prolific [^12] https://www.prolific.com/ from the target accent regions.
+To evaluate accent similarity between the generated speech and the source speech (Accent SMOS (src)) in accent-preserving VC, we randomly sample 20 utterances per selected DSRT, with speaker pairs drawn randomly from all 44 possible pairs (mapping from 11 American Midwest, Scottish, and Oceanian speakers to 4 Southern English speakers, covering both seen and unseen source accents).
+This results in 300 ratings per system per source accent region.
+To evaluate accent similarity between the generated speech and the target accent reference clip (Accent SMOS (tgt)) in accent-adaptive VC, we randomly sample 30 utterances per selected DSRT, again using random speaker pairs from the same 44 possible mappings.
+This results in 450 ratings per system.
+To evaluate speaker similarity between the generated speech and the target speaker reference clip (Speaker SMOS (tgt)) in both accent-preserving and accent-adaptive VC, we randomly sample 15 utterances per selected DSRT from the same pool of 44 speaker pairs, yielding 225 ratings per system.
+4.4 Evaluating Information Accessibility
+Information accessibility is evaluated only on the accents seen during HiFiGAN training, ensuring sufficient numbers of speakers for each accent.
+As described in Section 3.3, a training set is required to select high-frequency words and accent-discriminative combinations, while a separate test set is used for ABX evaluation to prevent speaker leakage.
+Accents outside the four seen accents contain no more than four speakers each, severely limiting the number of valid ABX triplets.
+We follow the same training–test partition as in HiFiGAN experiments for simplicity.
+Accent ABX is conducted using this partition, whereas phone and speaker ABX are computed only on the test split, as they do not require a training phase.
+Test experiments used all utterances in the split, except for the first 24 utterances to avoid identical sentences across all train and test speakers, which could bias the ABX scores toward lexical-context matching instead of reflecting true accent or speaker discrimination.
+Phone-level alignment (phone boundaries and labels) required by phone ABX, were obtained through the Montreal Forced Aligner with the English ARPABET dictionary [mfa_english_us_arpa_acoustic_2024, gorman2011prosodylab].
+We treat the accent ABX score computed on HuBERT layer 6 continuous features as an upper bound, as this layer exhibits the strongest accent-related information.
+Similarly, we use the phonetic and speaker ABX scores computed on HuBERT-ft layers 3 and 9 continuous features as upper bounds, respectively, since these layers demonstrate the highest phonetic and speaker recoverability in Section 5.
+In addition, we evaluate the accent ABX error rate of the accent-discriminative GenAID features, which achieves a low error rate of 9.87%, validating the effectiveness of the proposed metric.
+(a) Cross-accent VC evaluation results for information recoverability.
+(b) ABX evaluation results for information accessibility.
+Figure 2: Accent, speaker, and phonetic information in DSRTs obtained across different layers and continuous speech representations.
+All codebook sizes and code dimensions are set to 1024 for fair comparison.
+(a) Cross-accent VC evaluation results for information recoverability.
+(b) ABX evaluation results for information accessibility.
 Figure 3: Accent, speaker, and phonetic information in different DSRTs across codebook sizes.
-Full version of Figure 3.
+5 Results
+5.1 Layer Choice Matters for Accent
+Layer choice has a prominent impact on the recoverability and accessibility of information in HuBERT DSRTs, shown in Figure 2 (dark green curve).
+Notably, accent, speaker and phonetic information are distributed differently across layers.
+(1) Accent information recoverability, reflected by Accent COS-SIM w.r.t.
+source speaker, is most prominent in mid–early HuBERT layers (L6 & L9), and decreases in earlier or latter layers, indicating that accent cues emerge after low-level acoustic processing but are progressively abstracted away in higher layers.
+(2) Speaker information is most recoverable in early HuBERT layers (L3), indicated by Speaker COS-SIM to the source speaker, and decreases monotonically with depth.
+Overall, speaker information is limited, as cross-accent VC aims to generate voice similar to that of the target speaker with speaker ID conditioning.
+(3) Phonetic information, measured by PPG Distance to the source speech, is most complete in middle layers (L9 & L12).
+Both earlier and later layers show reduced phonetic completeness, reflecting a trade-off between raw acoustic detail and higher-level abstraction.
+(4) Information accessibility, measured by ABX error rates (see Figure 2(b)), reveals weak accent accessibility for HuBERT across layers, despite clear inter-layer variations.
+Accent accessibility peaks at L12, which differs from the recoverability peak at L6 or L9.
+This discrepancy highlights the necessity of resynthesis-based recoverability evaluation when assessing DSRTs for accent generation: while ABX accessibility captures relative trends, it cannot reliably identify optimal layers for synthesis.
+5.2 ASR Supervision Removes Accent Information
+ASR supervision reduces the recoverability and accessibility of accent information in DSRTs across layers, shown in Figure 2 (blue curve for HuBERT-ft, orange curve for Whisper).
+(1) HuBERT consistently shows higher accent recoverability and accessibility than HuBERT-ft or Whisper, particularly in latter layers, which are more affected by ASR supervision.
+(2) Whisper exhibits a peak in accent and phonetic information at intermediate layers (L12 & L15), different from the distribution of information in HuBERT or HuBERT-ft.
+Table 3: Objective evaluation of proposed content and content-accent tokens, compared with content and content-style tokens from Vevo [zhang2025vevo].
+Differences in DSRT design choices are marked in red, with best performances marked in bold.
+*: For content tokens, we want little accent information recovered from DSRTs, i.e. low A-SIM (src) and high A-SIM (tar). For content-accent or content-style tokens, we want the opposite.
+Token type
+Representation
+& Layer
+Codebook
+size
+A-SIM
+A-SIM
+S-SIM
+[…]
+S-SIM
+[…]
+[…]
+Vevo Content HuBERT L18 32 0.8143 0.9369 0.6817 0.9651 0.1718 0.0824
+Content-style HuBERT L18 8192 0.8923 0.8888 0.6989 0.9499 0.1427 0.0425
+Proposed Content HuBERT-ft L18 256 0.8081 0.9509 0.6782 0.9667 0.1488 0.0523
+Content-accent HuBERT L9 8192 0.9541 0.8227 0.7529 0.9104 0.1265 0.0344
+Table 4: Subjective evaluation of proposed content and content-accent tokens, compared with content and content-style tokens from Vevo [zhang2025vevo], with best performances marked in bold.
+95% confidence intervals are reported.
+[…]
+American Midwest Oceanian Scottish
+Vevo content-style 2.39±0.15 2.81±0.15 3.16±0.14 3.18±0.17
+Proposed content-accent 3.61±0.14 3.39±0.14 3.59±0.13 2.65±0.18
+[…]
+Southern English
+Vevo content 2.82±0.12 3.40±0.17
+Proposed content 3.15±0.12 3.60±0.16
+5.3 Reducing Codebook Size Provides Limited Disentanglement
+We then vary codebook sizes for three selected layers: HuBERT L18 (a common choice, as in Vevo [zhang2025vevo]), HuBERT-ft L18 (a layer with little accent information while retaining high phonetic information), and HuBERT L9 (a layer with most accent information), shown in Figure 3.
+(1) Focusing on accent recoverability, varying codebook sizes has a markedly smaller effect than changing representation models or layers.
+For each layer, increasing codebook sizes from 32 to 2048 consistently improves accent and phonetic recoverability across all inspected DSRTs.
+However, the recoverability gap induced by codebook sizes variation within a layer is substantially smaller than that induced by changing layers at a fixed codebook sizes.
+(2) Speaker and phonetic recoverability also increase with larger codebook sizes.
+Taking HuBERT DSRTs as an example, Accent COS-SIM drops sharply as codebook sizes decreases from 1024 to 32, while PPG distance also rises 20.4%.
+Despite speaker identity being partially suppressed by the injected speaker ID during resynthesis, speaker recoverability also declines.
+The parallel degradation across accent, speaker, and phonetic metrics suggests minimal disentanglement when reducing codebook size.
+The VQ bottleneck acts more as a lossy compressor for all information rather than a filter for specific features.
+5.4 Limitations of Commonly Used DSRTs in Speech Generation
+Based on the above findings, we highlight the limitations of commonly used DSRTs here.
+(1) Vevo’s choice of using DSTRs based on HuBERT L18 with codebook size 32 as content tokens and codebook size 8,192 as content-style tokens [zhang2025vevo] is suboptimal.
+Significant accent information is likely already lost and unrecoverable in HuBERT L18 - leading to insufficient accent information encoded in content-style tokens.
+Since naive reduction of codebook size cannot remove accent without harming phonetic information, content tokens achieve little accent information at the cost of intelligibility in resynthesised speech.
+(2) ASR-supervised DSRTs, adopted by [du2024cosyvoice], exhibit lower peak accent and phonetic recoverability than HuBERT DSRTs.
+Similar to content-style tokens, ASR-supervised DSRTs cannot encode accent sufficiently.
+5.5 Proposed Content and Content-Accent Tokens
+To enable better accent generation and control, we propose content-accent and content tokens for accent-preserving and accent-adaptive VC, respectively.
+A comparison between the proposed tokens and the tokens used in Vevo, together with the objective evaluation metrics, is presented in Table 3.
+The listening test results are reported in Table 4, which further validate the objective findings.
+Overall, the proposed content and content-accent tokens outperform Vevo's content and content-style tokens across most metrics, particularly in generating the desired accent while maintaining high content consistency and intelligibility.
+These results suggest improved disentanglement between accent and content.
+However, while the proposed content tokens achieve high speaker similarity to the target speaker in accent-adaptive VC, the proposed content-accent tokens perform worse in accent-preserving VC.
+This indicates that accent information remains partially entangled with speaker characteristics in the content-accent tokens.
+The lower speaker similarity observed for content-accent tokens may also be influenced by speaker–accent entanglement in both objective and subjective evaluations.
+As discussed in Section 4.3, the speaker embeddings used in the system encode some accent information.
+Consequently, when the generated speech more closely matches the source speaker's accent, the measured similarity to the target speaker may decrease.
+A similar phenomenon may occur in the subjective evaluation.
+Although listeners were instructed to ignore accent when rating speaker similarity, accent differences may still influence perceived speaker identity, as listeners often associate distinct accents with different speakers.
+This raises the broader question of whether accent and speaker characteristics can be fully disentangled in practice, particularly given that training data rarely contains the same speaker producing multiple accents.
+Additional supervision, such as explicit accent and speaker classification, may be required to further disentangle these factors.
+6 Discussion
+We observed a systematic delay of the optimal accessible layer relative to the recoverability peak.
+As discussed in [yeh2024estimating], earlier layers retain more complete acoustic information that can be effectively extracted by a strong HiFiGAN vocoder, whereas later layers gain phonetic discriminability from ASR-related objectives,
+Across both metrics, accent information follows a distinct yet intermediate trend between phonetic and speaker information.
+This aligns with linguistic intuition: accent is partially realised through systematic phonetic variation and partially through speaker-specific pronunciation patterns.
+Layer selection enables accent control precisely because accent, speaker, and phonetic information are distributed differently across layers.
+Consequently, a single "style layer", as suggested by the content–style token design in Vevo, is insufficient to jointly capture multiple speech attributes such as content, accent, and emotion.
+A more thorough investigation of more speech attributes, and how their information are represented across multiple layers and representations, is therefore necessary to support better controllable speech generation.
+Finally, our findings provide potential explanations for the hallucinated accents by ZS-TTS systems [zhong2025accentbox].
+Many ZS-TTS systems rely on representations (often deep or supervised layers) where accent recoverability is already attenuated, encouraging the model to guess accent patterns in inference or default to mainstream accents in training data.
+7 Conclusion
+In this work, we propose a framework for evaluating discrete speech representation tokens (DSRTs) from both recoverability and accessibility perspectives, in terms of accent, speaker, and phonetic information encoded, with a focus on providing guidance for accent generation and control.
+Based on the results using this framework on HuBERT- and Whipser-based DSRTs, we propose content and content-accent tokens, achieving more controllable accent in voice conversion (VC) than existing works.
+8 Acknowledgments
+[DEDACTED]
+9 Generative AI Use Disclosure
+During the preparation of this manuscript, the authors used ChatGPT (version 5.2 and 5.3) for language editing and readability improvement.
+The authors subsequently reviewed and revised the manuscript and take full responsibility for the final content.